OpenStudy (anonymous):

simplify complete! Please help i have the answer but i must understand how to get to the answer thank you! 24x+24 48x-16 --------- * --------- 12x-14 6x^2-6

OpenStudy (anonymous):

\[\frac{ 24x+24 }{ 12x-14 } * \frac{ 48x-16 }{ 6x^2-6 }\]

jhonyy9 (jhonyy9):

24x+24=24(x+1 12x-14=2(6x-7) 48x-16=16(3x-1) 6x^2 -6=6(x^2 -1)=6(x-1)(x+1) 24x+24 48x-16 24(x+1) 16(3x-1) 4*8*(3x-1) -------- * --------- = -------- * ----------- = ------------- = 12x-14 6x^2 - 6 2(6x-7) 6(x-1)(x+1) (6x-7)(x-1) 32(3x-1) = ------------ (6x-7)(x-1) hope this will help you sure easy understanding good luck bye